Derek left the ranch on horseback, riding at 21km/h. One hour later his brother Jonathon left the ranch to catch up with Derek. Jonathon rode at 25km/h. How long did it take for Jonathon to catch up to Derek?
Answer by checkley71(8403)

21T=25(T-1)
21T=25T-25
25T-21T=25
4T=25
T=25/4
T=6.25 HOURS AFTER DEREK LEFT THEY WILL MEET.
PROOF
21*6.25=25*5.25
131.25=131.25